# Gaussian-Process-based Adaptive Tracking Control with Dynamic Active Learning for Autonomous Ground Vehicles

### v1

@@ V-D Bayesian Optimization
$\quad{W}\in[-{w}_{\mathrm{b}}/2,w_{\mathrm{b}}/2],$ (28c)
@@
After solving (28) the synthesized reference is applied on the vehicle to obtain the new training dataset. Then, (28) can be solved again with the expanded dataset, achieving dynamic active learning.
@@ VI $\mathcal{L}_{2}$-gain Analysis
where ${f}_{\mathrm{cl}}:\mathcal{X}\times\mathcal{W}\times\Xi\rightarrow\mathcal{X}$ corresponds to the true vehicle dynamics (15) with (12d)-(12f) and (14) interconnected with the lateral and longitudinal controllers (23) and trained GP-based compensators (22). The argument $\xi$ describes the vector of uncertain model parameters, such as drivetrain parameters, cornering stiffnesses, vehicle mass and inertia parameters that describe different realizations of the true vehicle and are assumed to be bounded in a closed set $\Xi\subset\mathbb{R}^{n_{\xi}}$. The nominal model is represented by the parameter vector $\xi_{0}$ in this set, for which the lateral and longitudinal controllers (23) are designed based on (24) with weighting matrices $Q_{\mathrm{lo}}$, $Q_{\mathrm{la}}$, $R_{\mathrm{lo}}$, $R_{\mathrm{la}}$. When $\xi=\xi_{0}$, then from the adaptation point of view model there is no mismatch, so there is actually no need for GP augmentation and adaptive controller.
@@
To shift the equilibrium to the origin, the longitudinal velocity $v_{\xi}$ and velocity reference $v^{r}$ are centered: let $\tilde{v}_{\xi}=v_{\xi}-v^{\mathrm{r}}_{\mathrm{cent}}$ and $\tilde{v}^{\mathrm{r}}=v^{\mathrm{r}}-v^{\mathrm{r}}_{\mathrm{cent}}$, where $v^{\mathrm{r}}_{\mathrm{cent}}=(v^{\mathrm{r}}_{\mathrm{max}}+v^{\mathrm{r}}_{\mathrm{min}})/2$. With these new variables ${x}=[q\;e_{s}\;\theta_{\mathrm{e}}\;\tilde{v}_{\xi}\;v_{\eta}\;\omega]^{\top}\in\mathcal{X}$ is the state vector, $w=[\tilde{v}^{\mathrm{r}}\;c]^{\top}\in\mathcal{W}$ is the generalized disturbance signal and $z$ is the generalized performance signal that contains the tracking errors as $z=h(x,w)=[{\color[rgb]{0,0,0}\tilde{v}}_{\xi}-{\color[rgb]{0,0,0}\tilde{v}}^{\mathrm{r}}\;\;e_{s}]^{\top}\in\mathcal{Z}$. Furthermore, to indicate that $\tilde{v}^{\mathrm{r}}$ is only active in the low-frequency range, we augment the system with a first-order, strictly proper, low-pass filter. As a result, the state space is extended by one extra dimension, corresponding to the state of the filter i.e. $\tilde{x}=[q\;e_{s}\;\theta_{\mathrm{e}}\;v_{\eta}\;\tilde{v}_{\xi}\;\omega\;{\color[rgb]{0,0,0}\tilde{v}^{r}_{\mathrm{f}}}]^{\top}\in\tilde{\mathcal{X}}$. As the filter is strictly proper, the direct feed-through is eliminated from the output equation, i.e., $z=h(\tilde{x})=[\tilde{v}_{\xi}-{\color[rgb]{0,0,0}\tilde{v}^{r}_{\mathrm{f}}}\;\;e_{s}]^{\top}$.
@@
For choice of $\xi\in\Xi$, as a first step, an initial sparse GP model is fitted, via the VFE method to obtain the hyperparameters and $M$ inducing points, using a data set with $N$ samples obtained from the nonlinear vehicle model represented by $\xi$ under an input signal $u$ that is fixed among all trials. Then, $N_{\mathrm{act}}$ iterations of dynamic active learning in terms of experiment design and batch-wise adaptation of the sparse GP are accomplished according to Secs. II-C and V with the hyperparameters fixed among trials. This results in a learned compensator (22) that is added to the nominal control law to form the adapted closed-loop system.
@@
The calculation of the $\mathcal{L}_{2}$-gain of the resulting closed loop relies on the theory of dissipative dynamical systems [32]. For a given $\xi$, system (29) is said to be dissipative w.r.t. a quadratic supply $s(w,z)=\gamma^{2}w^{\top}w-z^{\top}z$ if there exists a non-negative storage function $V:\tilde{\mathcal{X}}\rightarrow\mathbb{R}^{+}$ such that, in case $V$ is differentiable, the differential dissipation inequality
@@
To overcome this difficulty, we propose an iterative, optimization-based approach, inspired by [33], which consists of two components: a learner and a verifier. First, the learner is responsible for finding a storage function candidate $V$ and corresponding $\mathcal{L}_{2}$-gain $\gamma$ by solving a convex optimization problem. To formulate the learner, we restrict ourselves to function candidates in the form of $V(x)=\tilde{x}^{\top}P(\tilde{x})\tilde{x}$, where $P(\tilde{x})=P_{0}+P_{1}\tilde{x}_{1}+\cdots+P_{7}\tilde{x}_{7}\succ 0$ as this function class remains linear in the parameters ${P}_{i=0}^{7}$, but is more flexible than the generic quadratic forms.
@@
where $\mathbb{X}$ and $\mathbb{W}$ are discrete grids, constructed by sampling the compact sets such that the sample points sufficiently cover $\tilde{\mathcal{X}}\times{\mathcal{W}}$. As (32c) is linear in the optimization variables, we can utilize this gridding approach even up to the case of 6 dimensions, as state-of-the-art numerical solvers can efficiently handle even a large number of linear constraints.
@@
As (33) is a small dimensional problem, numerical solvers with efficient algorithmic differentiation (e.g. CasADi [34] - IPOPT [35]) are capable of handling it. Note also that since (35) is nonlinear, there is no guarantee of finding the global optimum. To avoid getting stuck at a local maximum, the optimization can be performed multiple times starting from different initial values. For more sophisticated solutions that can provide mathematical guarantees to cover the entire search space, we can combine this simple heuristic approach with scenario method [36], or use adaptive sampling similar to that proposed in [37]. After solving (33), if $J$ is positive, the corresponding $\tilde{x}$ and $w$ values are added to the discrete sets $\mathbb{X}$, $\mathbb{W}$ and the iteration is repeated until either (32) becomes infeasible which means that we cannot obtain a bound for the induced $\mathcal{L}_{2}$-gain with the proposed storage function structure or the optimal value of $J$ remains negative which means that an upper bound for $\mathcal{L}_{2}$-gain is found with the storage function $V$, also showing stability.
@@ VII-A Setup and Simulation Environment
The performance of the proposed control architecture is first analyzed in a simulation environment. For the simulation study, we have developed a digital-twin model F1TENTH cars, corresponding to the dynamics discussed in Sec. III-B, using MuJoCo [38], a high-fidelity physics engine, which allows us to reliably test the adaptive schemes, as physical parameters can be easily altered. The nominal model corresponds to a parameter configuration of the MuJoCo model, where the parameters were obtained by identifying a real F1TENTH car, see [24]. The simulation environment and the model parameters are available on GitHub [^2] https://github/com/AIMotionLab-SZTAKI/AiMotionlab-Virtual.
To evaluate the performance of the adaptive controller approach, we artificially generated a large model mismatch in the digital-twin. The friction coefficients ($f_{\mathrm{wheel}}$) between the wheels and the ground have been reduced, while their radius ($r_{\mathrm{wheel}}$) and the overall inertia of the vehicle have been increased. The original and altered model parameters are displayed in Tab I. Furthermore, the steering dynamics have also been altered by scaling and adding an offset to it:
@@
For the synthesis of the nominal control laws (23) using the baseline single track model, the parameter-dependent gain matrices are obtained by solving (24), which has been implemented in Python with CVXPY [39] and solved with Mosek [^3] https://www.mosek.com. The weighting matrices of the LQR have been tuned using numerical simulations with the model resulting in $Q_{\mathrm{la}}=\mathrm{diag}(1,80,0)$, $R_{\mathrm{la}}=500$ and $Q_{\mathrm{lo}}=1$, $R_{\mathrm{lo}}=100$ for the lateral and longitudinal controller, respectively. Furthermore, the gain of the virtual velocity reference generator is $k_{v}=0.1$.
@@ VII-B Data Collection and GP Training
For initial GP training, we used lemniscate trajectories (Fig. 4) with constant speeds (0.75-1.25-2 m/s) and 25 Hz sampling, resulting in a dataset of $N={12000}$ points. The initial training of the hyperparameters of the GPs has been performed offline by maximizing the VFE cost (8) by gradient descent with $M=30$. Next, we used the dynamic active learning method from Sec. V to refine the GP components. The algorithm used a lemniscate initial trajectory and a 1.25 m/s constant speed profile were used, with $w_{\mathrm{b}}=0.5$, $v_{\mathrm{min}}=0.5$ m/s, and $v_{\mathrm{max}}=2$ m/s as bounds. We have used $n=7$ equidistant nodes along the path, yielding 14 optimization variables. variables were optimized using 25 initial samples and 20 iterations per trajectory synthesis. The optimization has been performed using 25 initial samples and a total of 20 iterations. The trajectories are shown in Fig. 4.
@@
Using the data collected along the trajectories, we retrained the GP components in a batch-wise manner with the augmented datasets. We performed 5 iterations of the dynamic active learning, where during each iteration, we collected additional training data along a new trajectory and retrained the GPs. At each iteration, for the sake of the analysis, we evaluated the cumulative variance (27) along a separate test trajectory to quantify how the active learning method improves the uncertainty of the approximation. As the results show in Fig. 5, the proposed dynamic active learning efficiently reduced the cumulative variance after each iteration.
@@ VII-C Adaptive Control
To demonstrate the efficiency of the proposed control algorithm, we compare three scenarios, where the vehicle tracks the same predefined trajectory. In the first scenario (LPV-LQ), only the nominal feedback laws (23) are used. Then, in the second and third scenarios, we utilize the GP-based adaptive compensator (22) along with the nominal controller. For online adaptation of the GPs, we start from the offline estimated GP and apply the RLS-GP method of [11] and the RGB-GP method of Sec. II-C with $Z=20$ batch size. Furthermore, for the RLS-GP $\gamma=0.995$ forgetting factor and $\beta=0.9$ confidence level has been used, while for the RGB-GP $n_{\alpha}=5$ update iteration has been selected to maintain computational feasibility, with $\alpha=0.1$ learning rate.
The simulation results are depicted in Fig. 6. As shown, due to the significant model mismatch, the nominal controller cannot track the reference accurately, as both the lateral ($e_{s}$), and the heading ($\theta_{\mathrm{e}}$) errors are significant. Furthermore, a steady-state error occurs in the longitudinal velocity ($v_{\mathrm{err}}$) that results in increased longitudinal position error ($s_{\mathrm{err}}$). On the other hand, we can observe that the adaptive controllers can efficiently decrease the tracking errors and ensure the precise tracking of the reference trajectories. A comparison between the control inputs generated by the LPV-LQ and the two GP-based adaptive algorithms is also displayed. Finally, to quantify the tracking performance, the maximum and root mean square (RMS) values of the tracking errors are collected in Tab. II. As the table and the figure show, the RBG-GP provides less noisy estimates and superior tracking performance compared to the RLS-GP.
TABLE II: Tracking errors of the nominal, the RLS-based (RLS-GP) and the RGB-based (RGB-GP) online adaptive controllers.
max($e_{s}$) max$(s_{\mathrm{err}})$ $\|e_{s}\|_{\mathrm{RMS}}$ $\|s_{\mathrm{err}}\|_{\mathrm{RMS}}$
LPV-LQ 0.12 1.16 0.07 0.83
RLS-GP 0.05 0.37 0.01 0.15
RBG-GP 0.04 0.28 0.01 0.14
Figure 6: Comparison of the simulation results using the nominal and the adaptive tracking controllers.

### v2

@@ V-D Bayesian Optimization
$\quad{W}\in[-{{\color[rgb]{0,0,0}\mathpzc{w}}}_{\mathrm{b}}/2,{{\color[rgb]{0,0,0}\mathpzc{w}}}_{\mathrm{b}}/2],$ (28c)
@@
After solving (28), the synthesized reference is applied on the vehicle to obtain the new training dataset. Then, (28) can be solved again with the expanded dataset, achieving dynamic active learning.
@@ VI $\mathcal{L}_{2}$-gain Analysis
where ${f}_{\mathrm{cl}}:\mathcal{X}\times\mathcal{W}\times\Xi\rightarrow\mathcal{X}$ corresponds to the true vehicle dynamics (15) with (12d)-(12f) and (14) interconnected with the lateral and longitudinal controllers (23) and trained GP-based compensators (22). The argument $\xi$ describes the vector of uncertain model parameters, such as drivetrain parameters, cornering stiffnesses, vehicle mass and inertia parameters that describe different realizations of the true vehicle and are assumed to be bounded in a closed set $\Xi\subset\mathbb{R}^{n_{\xi}}$. The nominal model is represented by the parameter vector $\xi_{0}$ in this set, for which the lateral and longitudinal controllers (23) are designed based on (24) with weighting matrices $Q_{\mathrm{lo}}$, $Q_{\mathrm{la}}$, $R_{\mathrm{lo}}$, $R_{\mathrm{la}}$. When $\xi=\xi_{0}$, then, from the adaptation point of view, there is no model mismatch, so there is actually no need for GP augmentation and adaptive controller.
@@
To shift the equilibrium to the origin, the longitudinal velocity $v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$ and velocity reference $v^{\mathrm{{r}}}$ are centered: let $\tilde{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}=v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-v^{\mathrm{r}}_{\mathrm{cent}}$ and $\tilde{v}^{\mathrm{r}}=v^{\mathrm{r}}-v^{\mathrm{r}}_{\mathrm{cent}}$, where $v^{\mathrm{r}}_{\mathrm{cent}}=(v^{\mathrm{r}}_{\mathrm{max}}+v^{\mathrm{r}}_{\mathrm{min}})/2$. With these new variables, ${x}=[q\;e_{s}\;\theta_{\mathrm{e}}\;\tilde{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\;v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\;\omega]^{\top}\in\mathcal{X}$ is the state vector, $w=[\tilde{v}^{\mathrm{r}}\;c]^{\top}\in\mathcal{W}$ is the generalized disturbance signal and $z$ is the generalized performance signal that contains the tracking errors as $z=h(x,w)=[{\color[rgb]{0,0,0}\tilde{v}}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-{\color[rgb]{0,0,0}\tilde{v}}^{\mathrm{r}}\;\;e_{s}]^{\top}\in\mathcal{Z}$. Furthermore, to indicate that $\tilde{v}^{\mathrm{r}}$ is only active in the low-frequency range, we augment the system with a first-order, strictly proper, low-pass filter. As a result, the state space is extended by one extra dimension, corresponding to the state of the filter i.e. $\tilde{x}=[q\;e_{s}\;\theta_{\mathrm{e}}\;v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\;\tilde{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\;\omega\;{\color[rgb]{0,0,0}\tilde{v}^{\mathrm{r}}_{\mathrm{f}}}]^{\top}\in\tilde{\mathcal{X}}$. As the filter is strictly proper, the direct feed-through is eliminated from the output equation, i.e., $z=h(\tilde{x})=[\tilde{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-{\color[rgb]{0,0,0}\tilde{v}^{r}_{\mathrm{f}}}\;\;e_{s}]^{\top}$.
@@
Given a choice of $\xi\in\Xi$, first an initial sparse GP model is fitted, via the VFE method to obtain the hyperparameters and $M$ inducing points, using a data set with $N$ samples obtained from the nonlinear vehicle model represented by $\xi$ under an input signal $u$ (fixed for all choices of $\xi\in\Xi$). Then, $N_{\mathrm{act}}$ iterations of dynamic active learning in terms of experiment design and batch-wise adaptation of the sparse GP are accomplished according to Secs. II-C and V with the hyperparameters fixed among trials. This results in a learned compensator (22) that is added to the nominal control law to form the adapted closed-loop system.
@@
The calculation of the $\mathcal{L}_{2}$-gain of the resulting closed loop relies on the theory of dissipative dynamical systems [34]. For the chosen $\xi$, the system (29) is said to be dissipative w.r.t. a quadratic supply $s(w,z)=\gamma^{2}w^{\top}w-z^{\top}z$ if there exists a non-negative storage function $V:\tilde{\mathcal{X}}\rightarrow\mathbb{R}^{+}$ such that, in case $V$ is differentiable, the differential dissipation inequality
@@
To overcome this difficulty, we propose an iterative, optimization-based approach, inspired by [19], which consists of two components: a learner and a verifier. First, the learner is responsible for finding a storage function candidate $V$ and corresponding $\mathcal{L}_{2}$-gain $\gamma$ by solving a convex optimization problem. To formulate the learner, we restrict ourselves to function candidates in the form of $V(x)=\tilde{x}^{\top}P(\tilde{x})\tilde{x}$, where $P(\tilde{x})=P_{0}+P_{1}\tilde{x}_{1}+\cdots+P_{7}\tilde{x}_{7}\succ 0$ as this function class remains linear in the parameters $\{P_{i}\}_{i=0}^{7}\in\mathbb{R}^{7\times 7}$, but is more flexible than the generic quadratic forms.
@@
where $\mathbb{X}$ and $\mathbb{W}$ are discrete grids, constructed by sampling the compact sets such that the sample points sufficiently cover $\tilde{\mathcal{X}}\times{\mathcal{W}}$. Note that the denser the generated grid, the more likely it is that the learner can synthesize a valid storage function without any additional sampling. As (32c) is linear in the optimization variables, we can utilize this gridding approach even up to the case of 6 dimensions, as state-of-the-art numerical solvers can efficiently handle even a large number of linear constraints.
@@
As (33) is a small dimensional problem, numerical solvers with efficient algorithmic differentiation (e.g. CasADi [35] - IPOPT [36]) are capable of handling it. Note also that since (35) is nonlinear, there is no guarantee of finding the global optimum. To avoid getting stuck at a local maximum, as a workaround, the optimization is performed multiple times starting from different initial values. For more sophisticated solutions that can provide mathematical guarantees to cover the entire search space, we can combine this simple heuristic approach with scenario method [37], or use adaptive sampling similar to that proposed in [38]. After solving (33), if $J$ is positive, the corresponding $\tilde{x}$ and $w$ values are added to the discrete sets $\mathbb{X}$ and $\mathbb{W}$. The iteration is then repeated until one of two conditions is met: (i) (32) becomes infeasible, indicating that the proposed storage function structure cannot provide a bound for the induced $\mathcal{L}_{2}$-gain; or (ii) the optimal value of $J$ remains negative, implying that an upper bound for the $\mathcal{L}_{2}$-gain and storage function $V$ has been found, which also demonstrates the stability of the system.
@@ VII-A Setup and Simulation Environment
The performance of the proposed control architecture is first analyzed in a simulation environment. For the simulation study, we have developed a digital-twin model F1TENTH cars, corresponding to the dynamics discussed in Sec. III-B, using MuJoCo [39], a high-fidelity physics engine, which allows us to reliably test the adaptive schemes, as physical parameters can be easily altered. The nominal model corresponds to a parameter configuration of the MuJoCo model, where the parameters were obtained by identifying a real F1TENTH car, see [25]. The simulation environment and the model parameters are available on GitHub [^2] https://github.com/AIMotionLab-SZTAKI/AiMotionlab-Virtual.
To evaluate the performance of the adaptive controller approach, we artificially generate a large model mismatch in the digital-twin. The friction coefficients ($f_{\mathrm{wheel}}$) between the wheels and the ground are reduced, while their radius ($r_{\mathrm{wheel}}$) and the overall inertia of the vehicle are increased. The original and altered model parameters are displayed in Tab I. Furthermore, the steering dynamics are also altered by scaling and adding an offset to it:
@@
For the synthesis of the nominal control laws (23) using the baseline single track model, the parameter-dependent gain matrices are obtained by solving (24), which has been implemented in Python with CVXPY [40] and solved with Mosek [^3] https://www.mosek.com. The weighting matrices of the LQR are tuned using numerical simulations with the model resulting in $Q_{\mathrm{la}}=\mathrm{diag}(1,80,0)$, $R_{\mathrm{la}}=500$ and $Q_{\mathrm{lo}}=1$, $R_{\mathrm{lo}}=100$ for the lateral and longitudinal controller, respectively. Furthermore, the gain of the virtual velocity reference generator is $k_{v}=0.1$.
@@ VII-B Data Collection and GP Training
For initial GP training, we use lemniscate trajectories (Fig. 4) with constant speeds (0.75-1.25-2 m/s) and 25 Hz sampling, resulting in a dataset of N=12000N={12000} points. The initial training of the hyperparameters of the GPs is performed offline by maximizing the VFE cost (8) by gradient descent with $M=30$. Next, we use the dynamic active learning method from Sec. V to refine the GP components. The algorithm uses a lemniscate initial trajectory and a 1.25 m/s constant speed profile is used, with ${{\color[rgb]{0,0,0}\mathpzc{w}}}_{\mathrm{b}}=0.5$, $v_{\mathrm{min}}=0.5$ m/s, and $v_{\mathrm{max}}=2$ m/s as bounds. We use $n=7$ equidistant nodes along the path, yielding 14 optimization variables. These variables are optimized using 25 initial samples and 20 iterations per trajectory synthesis. The optimization is performed using 25 initial samples and a total of 20 iterations. The trajectories are shown in Fig. 4.
@@
Using the data collected along the trajectories, we retrain the GP components in a batch-wise manner with the augmented datasets. We perform 5 iterations of the dynamic active learning, where during each iteration, we collect additional training data along a new trajectory and retrain the GPs. At each iteration, for the sake of the analysis, we evaluate the cumulative variance (27) along a separate test trajectory to quantify how the active learning method improves the uncertainty of the approximation. As the results show in Fig. 5, the proposed dynamic active learning efficiently reduces the cumulative variance after each iteration.
@@ VII-C Adaptive Control
To demonstrate the efficiency of the proposed control algorithm, we compare four scenarios, where the vehicle tracks the same predefined trajectory. In the first scenario (LPV-LQ), only the nominal feedback laws (23) are used. Then, in the second and third scenarios, we utilize the GP-based adaptive compensator (22) along with the nominal controller. Finally, in the fourth scenario, we employ the model predictive contouring control with a GP-augmented [^4] Utilizing the MPCC algorithm without the GP-augmentation resulted in infeasibilities due to the large model mismatch. dynamic single track model (MPCC-GP) from [9] for comparison. However, it is important to note that the MPCC formulation aims to maximize progress along the path instead of reference trajectory tracking, therefore, for the analysis, longitudinal error measurements are omitted for this scenario. For online adaptation of the GPs, we start from the offline estimated GP and apply the RLS-GP method of [11] and the RGB-GP method of Sec. II-C with $Z=20$ batch size. Furthermore, for the RLS-GP $\gamma=0.995$ forgetting factor and $\beta=0.9$ confidence level are used, while for the RGB-GP $n_{\alpha}=5$ update iterations are selected to maintain computational feasibility, with $\alpha=0.1$ learning rate. To retain real-time feasibility, we used real-time iteration (RTI) for the MPCC scheme.
The simulation results are represented in Fig. 6 and Tab. II. As shown, due to the significant model mismatch, the nominal controller cannot track the reference accurately, as both the lateral ($e_{s}$), and the heading ($\theta_{\mathrm{e}}$) errors are significant. Furthermore, a steady-state error occurs in the longitudinal velocity ($v_{\mathrm{err}}$) that results in increased longitudinal position error ($s_{\mathrm{err}}$). On the other hand, we can observe that the adaptive controllers can efficiently decrease the tracking errors and ensure the precise tracking of the reference trajectories. A comparison between the control inputs generated by the LPV-LQ and the two GP-based adaptive algorithms is also displayed. Finally, to quantify the tracking performance, the maximum and root mean square (RMS) values of the tracking errors are collected in Tab. II. As the table and the figure show, the RBG-GP provides less noisy estimates and superior tracking performance compared to the RLS-GP. Furthermore, it can also be seen that the method has similar performance to the MPCC algorithm for a fraction of the computational complexity. This is partially due to the simplicity of the LPV-LQ compared to the MPCC, which requires the solution of an optimization problem at each iteration. Moreover, when dealing with GP-augmented models, the GPs have to be evaluated along the whole control horizon of the MPCC, whereas our proposed solution only requires one evaluation at each control cycle. We furthermore note that, due to the utilization of sparse GPs, the initial offline training data size does not affect the real-time performance of the method.
TABLE II: Tracking errors (in meters) of the nominal, the RLS-based (RLS-GP) and the RGB-based (RGB-GP) online adaptive controllers, and average control cycle time (in milliseconds).
[…]
LPV-LQ 0.12 1.16 0.07 0.83 0.02
RLS-GP 0.05 0.37 0.01 0.15 0.77
RBG-GP 0.04 0.28 0.01 0.14 4.27
GP-MPCC 0.06 0.02 21.14
Figure 6: Comparison of the simulation results using the nominal and the adaptive tracking controllers as well as the MPCC as a baseline.